OpenStudy (anonymous):

if price of CNG increases by 20% and sara intend to spend only 5% more on CNG by how much % should she reduce the quantity of CNG she buys??

OpenStudy (kropot72):

Let a1 be the amount that would have been bought at the original price, and let a2 be the reduced amount that is required at the new price to meet the spending target. \[\large 1.2a _{2}=1.05a _{1}\] \[\large \frac{a _{2}}{a _{1}}=\frac{1.05}{1.2}=0.875\] Therefore the required percentage reduction is given by: \[\large (1-0.875) \times100=(you\ can\ calculate)\ percent\]

OpenStudy (anonymous):

ty @kropot72

OpenStudy (anonymous):

u done it wrong @kropot72

OpenStudy (anonymous):

@Haseeb96

OpenStudy (kropot72):

@rosebird. I double checked and my method is correct. What value of percentage did you calculate using my method?

OpenStudy (haseeb96):

@rosebird what are the choices written in your questions book? And however the method used above is correct according to me.

OpenStudy (haseeb96):

And the answer is also correct as the above man done that. I think the the quantity of CNG will be reduced is 12.5% by her

OpenStudy (anonymous):

ok @Haseeb96 @kropot72

OpenStudy (anonymous):

yeh its correct now i am getting the write answer @kropot72

OpenStudy (kropot72):

Cool! You're welcome :)

OpenStudy (anonymous):

:) @kropot72